# Bounding the conditional von-Neumann entropy for device independent cryptography and randomness extraction

### v1

@@ III Applications and numerical examples
In this section, we present numerical evidence supporting the methods and tools developed in section II. To structure this discussion, we divide section III into two parts, as outlined in [5].
In the first part, we examine a CHSH game scenario with two possible inputs and two possible outputs. Our objective here is to compute $H(A|X=0,Q_{E})$, which is a quantity of particular interest in two main areas: one-sided device-independent randomness extraction (DIRE) and device-independent quantum key distribution (DIQKD). In one-sided DIRE, $H(A|X=0,Q_{E})$ provides a measure of the randomness that can be reliably extracted from Alice’s outcomes, assuming a fixed input $X=0$ and some quantum side information $Q_{E}$. Similarly, in DIQKD, this conditional entropy helps in determining the amount of secure randomness extractable from the quantum correlations between Alice and an adversary.
@@
For clarity, we abbreviate scenarios based on the number of inputs and outputs per party. Specifically, a scenario with $n_{a}$ inputs for Alice, $n_{b}$ inputs for Bob, $o_{a}$ possible outcomes for each input of Alice, and $o_{b}$ possible outcomes for each input of Bob is denoted by $n_{a}n_{b}o_{a}o_{b}$. For instance, the CHSH scenario with two inputs and two outputs per party is a $2$-scenario.
@@
This quantity remains significant even in finite-sample-size regimes due to recent advancements in entropy accumulation theorems (EAT) [19, 20, 21, 22, 23]. These theorems provide a framework for quantifying entropy in finite scenarios by demonstrating that finite-size quantities, like the smoothed min-entropy, can be bounded from below by a sum of asymptotic quantities. Thus, an essential part of a DIQKD security proof reduces to reliably lower bounding the asymptotic quantities in (17).
Beyond DIQKD, (16) is also important for randomness extraction, a process aimed at generating bits that are random and independent of any other system from a quantum experiment [24, 25, 26, 41].
As a concrete example of one-sided randomness extraction, consider the $2$-scenario under the constraint that the underlying, potentially unknown, state $\rho$ violates the CHSH inequality (Clauser-Horne-Shimony-Holt) with a value of $\omega\in[2,2\sqrt{2}]$:
@@
Notably, in $2$-scenarios, the local polytope is uniquely described by the CHSH inequality alone, which is sufficient to certify non-locality in such settings [31]. Furthermore, in this specific $2$-scenario, with only the CHSH inequality as a constraint, an analytic solution for (16) is achievable in a device-independent context, as demonstrated by [9]. Therefore, we use this as a benchmark in Figure 1.
For the experiment shown in Figure 1 in a $2$-scenario with a CHSH inequality as the constraint, we employ a second-level NPA hierarchy with additional monomials of the form $M_{a|x}N_{b|y}P^{(a)}$. We also apply the speedups for numerical evaluations discussed in Appendix E. This modification yields a valid lower bound that serves as a good approximation. In conclusion, the optimization approach is highly efficient, allowing us to partition the interval into many subintervals while keeping the runtime manageable, even on standard personal computers, with computation times remaining within a few seconds.
@@
This expression represents the conditional entropy of the joint outcomes $A$ and $B$, given fixed measurement settings $X=\tilde{x}$ and $Y=\tilde{y}$, as well as quantum side information $Q_{E}$. This joint entropy quantifies the total randomness that can be extracted from the outcomes of both parties under given conditions, making it particularly relevant for applications that require high degrees of randomness.
@@
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E}).$ (20)
@@
First, we generalize the CHSH-game presented in subsection III.1 to a two-sided setting. In this case, the setup is governed by the CHSH inequality (18), and we examine a $2$-scenario where both Alice and Bob have two measurement inputs, each yielding one of two possible outcomes. The resulting conditional entropy values are similar to those obtained in [3, Fig. 2] and are visualized in Figure 1. To establish upper bounds in this setting, we formulate an optimization problem as described in [5], enabling us to assess lower bounds on randomness extractable under CHSH constraints.
@@ 2 $\pi/2$ $\pi/8$
We show in Figure 1 that we match at least the upper bound with $2$ random bits. Moreover, comparing this result with [3, Fig. 3] questions whether a third basis on Bob’s side has an advantage.
IV Summary
@@
This work focuses on optimizing the conditional von-Neumann entropy within a device-independent framework. From a broader perspective, we ask how much randomness can be extracted from two spatially separated, hypothetical “black boxes” located in Alice’s and Bob’s respective labs, where the outcomes follow the statistical laws of quantum theory. Instead of specifying a concrete quantum experiment represented by a Hilbert space $\mathcal{H}$ and a trace-class operator $\rho$ on $\mathcal{H}$, we rely on the achievable statistics defined by general quantum theory. The core principle underlying the presence of genuine randomness here is the monogamy of correlations, which restricts information sharing in non-signalling theories as demonstrated in [42, 43]. This principle remains applicable even with only limited knowledge of the output statistics in (2). Additionally, bounding the capabilities of Alice, Bob, and Eve to the framework of quantum theory is significant not only for applications but also for advancing our understanding of quantum principles. For instance, [44] exploits the gap between general non-signalling theories and quantum theory, highlighting this difference as a powerful tool.
@@
Our approach serves as a versatile framework for deriving lower bounds on the conditional von-Neumann entropy. The main technical contributions of our work include the application of the integral representation from [27] adapted for this setting, as well as the formulation of a non-commutative polynomial optimization problem, which we solve using the NPA hierarchy [37, 35]. This approach bridges the gap between two established methods: those in [3] and [4]. The technique from [3] has become standard in security proofs for device-independent quantum key distribution (DIQKD) [23] and has since been extended to Petz-Rényi divergences [45]. Compared to [3], our method requires fewer NPA variables, as our projections $P_{k}^{(a)}$ in Theorem II.2 reduce complexity relative to the non-Hermitian operators $Z_{a,i}$ in [3]. This reduction streamlines the operator set in the sum-of-squares cone and enables us to impose additional constraints within the kernel of the map $\varphi$ in [36, Fundamental Lemma]. Moreover, our method builds on discretization techniques for integration from [32], while [3] relies on the Gauss-Radau quadrature method from numerical integration theory. However, the advantages of each integration method for the specific computations required remain open for investigation. The method in [4], though distinct from these approaches, recasts the conditional von-Neumann entropy in (1) as
@@
interpreting it as an entropy production. Here, the channel $T_{Q_{A}\to A}$ acts as the measurement channel on Alice’s system. The method in [4] establishes lower bounds through a variational approach, employing the Gibbs variational principle and generalized Golden-Thompson inequalities via multivariate trace inequalities [46]. A direct comparison of these conceptually related integral-based methods with the entropy production approach would be a valuable avenue for future research. In section III, we present multiple scenarios to illustrate the effectiveness of our method for randomness extraction. These scenarios include both numerical validations and theoretical comparisons to benchmark results from existing studies.
@@
Firstly, we analyze a $2222$ scenario (two measurement settings per party, each with two outcomes) under CHSH constraints, which is a standard setting in quantum information theory for testing Bell inequalities. We consider this scenario for both one-sided and two-sided randomness extraction, where randomness is certified by outcomes that are either private to one party (one-sided) or to both parties (two-sided). Our results in this $2222$ scenario show strong agreement with the analytical values reported in [9], demonstrating the accuracy and robustness of our approach. Additionally, we explore a more intricate $2322$ scenario, where one party has three measurement settings while the other has two. This scenario is used to examine global randomness extraction, which involves jointly certifying randomness across all outcomes. Our findings, obtained under the statistical framework of Werner states, indicate that this more complex setup does not offer a substantial qualitative advantage over the $2222$ scenario. This suggests that increasing the measurement settings in this configuration may not necessarily enhance randomness extraction, a result that merits further investigation. An intriguing direction for future research would be to systematically compare different configurations - such as the $2222$, $2322$ and $3322$ scenarios (where the latter allows each party three measurement settings with two outcomes) - to determine how these variations impact randomness generation under different assumptions of honest implementation or even a random key basis [30]. This could uncover new insights into the optimal design of randomness extraction protocols. Moreover, it would be valuable to examine randomness extraction in more complex non-local games, such as the magic square game discussed in [44]. These games offer alternative frameworks for randomness certification that might provide unique advantages in terms of security and efficiency in quantum protocols.
V Acknowledgements
GK thanks Mario Berta for very helpful advice around the topic. GK thanks Marco Tomamichel for hosting him at CQT Singapore in fall 2024 and Ian George for many helpful discussions around (DI)QKD. The authors thank Omar Fawzi for a discussion around the topic on an early stage of the project.

### v2

@@ III Applications and numerical examples
In this section, we demonstrate the methods and tools developed in section II. To structure this discussion, we divide section III into two parts:
In the first part, we examine a CHSH game scenario with two possible inputs and two possible outputs. Our objective here is to compute $H(A|X=0,Q_{E})$, which is a quantity of particular interest in two main areas: one-sided device-independent randomness extraction (DIRE) and device-independent quantum key distribution (DIQKD). In one-sided DIRE, $H(A|X=0,Q_{E})$ provides a measure of the randomness that can be reliably extracted from Alice’s outcomes, assuming a fixed input $X=0$ and some quantum side information $Q_{E}$. Similarly, in DIQKD, this conditional entropy helps in determining the asymptotic amount of secure randomness extractable from the quantum correlations between Alice and an adversary.
@@
For clarity, we abbreviate scenarios based on the number of inputs and outputs per party. Specifically, a scenario with $n_{a}$ inputs for Alice, $n_{b}$ inputs for Bob, $o_{a}$ possible outcomes for each input of Alice, and $o_{b}$ possible outcomes for each input of Bob is denoted by $n_{a}n_{b}o_{a}o_{b}$. For instance, the CHSH scenario with two inputs and two outputs per party is a 2222-scenario. Furthermore we observe that all quantities under consideration in this work are usable for spot-checking protocols. An interesting direction for future research is to extend this to averaging the entropy terms without spot-checking.
@@
This quantity remains significant even in finite-sample-size regimes due to recent advancements in entropy accumulation theorems (EAT) [13, 14]. These theorems provide a framework for quantifying entropy in finite scenarios by demonstrating that finite-size quantities, like the smoothed min-entropy, can be bounded from below by a sum of asymptotic quantities. Thus, an essential part of a DIQKD security proof reduces to reliably lower bounding the asymptotic quantities in (18).
Beyond DIQKD, (17) is also important for randomness extraction, a process aimed at generating bits that are random and independent of any other system from a quantum experiment [15, 16, 17].
As a concrete example of one-sided randomness extraction, consider the 2222-scenario under the constraint that the underlying, potentially unknown, state $\rho$ violates the CHSH inequality (Clauser-Horne-Shimony-Holt) with a value of $\omega\in[2,2\sqrt{2}]$:
@@
Notably, in 2222-scenarios, the local polytope is uniquely described by the CHSH inequality alone, which is sufficient to certify non-locality in such settings [27]. Furthermore, in this specific 2222-scenario, with only the CHSH inequality as a constraint, an analytic solution for (17) is achievable in a device-independent context, as demonstrated by [34]. Therefore, we use this as a benchmark in Figure 2.
For the experiment shown in Figure 2 in a 2222-scenario with a CHSH inequality as the constraint, we employ a second-level NPA hierarchy with additional monomials of the form $M_{a|x}N_{b|y}P^{(a)}$. We also apply the speedups for numerical evaluations discussed in Appendix E. This modification yields a valid lower bound that serves as a good approximation. In conclusion, the optimization approach is highly efficient, allowing us to partition the interval into many subintervals while keeping the runtime manageable, even on standard personal computers, with computation times remaining within a few seconds.
@@
This expression represents the conditional entropy of the joint outcomes $A$ and $B$, given fixed measurement settings $X=\tilde{x}$ and $Y=\tilde{y}$, as well as quantum side information $Q_{E}$. This joint entropy quantifies the total randomness that can be extracted from the outcomes of both parties under given conditions, making it particularly relevant for applications such as random number generator.
@@
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E})_{\rho_{ABQ_{E}|\tilde{x},\tilde{y}}}$
@@
First, we generalize the CHSH-game presented in subsection III.1 to a two-sided setting. In this case, the setup is governed by the CHSH inequality (19), and we examine a 2222-scenario where both Alice and Bob have two measurement inputs, each yielding one of two possible outcomes. The resulting conditional entropy values are similar to those obtained in [22, Fig. 2] and are visualized in Figure 2. To establish upper bounds in this setting, we formulate an optimization problem as described in [32], enabling us to assess lower bounds on randomness asymptotically extractable under CHSH constraints.
@@ 2 $\pi/2$ $\pi/8$
We show in Figure 2 that we match at least the upper bound with $2$ random bits. Moreover, comparing this result with [22, Fig. 3] questions whether a third basis on Bob’s side has an advantage.
IV Beyond CHSH
CHSH has long been the primary Bell inequality used for device-independent randomness. This is mainly because the $2222$-scenarios enjoy the powerful advantage that the $C^{\star}$-algebra of two projections is well-understood and is essentially as intricate as qubit-based information theory. Many seminal results have been obtained by reducing the device-independent scenario on the algebra of two projections back to the qubit case [11].
However, one can easily conceive of protocols featuring significantly different arrangements of input-output scenarios, which then naturally give rise to different families of Bell inequalities whenever quantum nonlocality arises. In such cases, a straightforward reduction to qubits or even qudits is often not feasible, necessitating either the use of self-testing methods for specific distributions, states and measurements or the acceptance of the full generality of solving the optimization in Theorem II.2.
In a recent result [24], DIQKD protocols beyond qubits were investigated, proposing key rates for more general scenarios than the CHSH and $2222$-cases. However, computing bounds on these key rates using the Brown–Fawzi–Fawzi (BFF) method can be challenging with standard computational resources. For this reason, the results in [24] were obtained by evaluating the min-entropy (i.e., by optimizing the guessing probability). With stronger computational resources, we present a comparison of our method with BFF for I3322 in Figure 4, and with BFF for CGLMP in Figure 6 and show that our method is on a comparable level faster and gives better bounds. Since our focus here is on the fundamental theoretical building blocks of entropy estimation in device-independent scenarios, we concentrate on calculating lower bounds on the asymptotically extractable randomness but in terms of bounds which are converging towards the conditional von Neumann entropy. As we will show numerically in the following, our tools are able to estimate the randomness for the CGLMP 2233-Bell inequality [35] and the I3322 Bell-inequality [36] significantly beyond the bounds from the min-entropy.
$p_{\operatorname{guess}}$ $(A|E)=$ (25)
$\sup$ $\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|\tilde{x}}C_{a}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},C_{a^{\prime}}]=[N_{b|y},C_{a^{\prime}}]=0$
$\quad b\in B,\ a,a^{\prime}\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y}$
$\sum_{a}C_{a}=\mathds{1},\ C_{a}\geq 0,\quad a\in A$
and the min entropy is then defined as
$H_{\operatorname{min}}(A|E)\coloneqq-\log p_{\operatorname{guess}}(A|E).$ (26)
IV.1 Randomness from CGLMP
Our first example concerns the CGLMP Bell inequality for a $2233$ scenario [35, eqs. (4) and (5)]. This is a Bell inequality for two parties, each having two measurement settings and three outcomes per setting. If we define, for generally correlated probability distributions, the expression
$P(A_{a}=B_{b}+k)\coloneqq\sum_{j=0}^{d-1}P\bigl{(}A_{a}=j,\;B_{b}=(j+k)\bmod d\bigr{)},$ (27)
then the CGLMP Bell inequality can be stated as
$I_{\operatorname{CGLMP}}\coloneqq$ $\Bigl{[}P(A_{1}=B_{1})+P(B_{1}=A_{2}+1)+P(A_{2}=B_{2})$ (28)
$\quad+P(B_{2}=A_{1})\Bigr{]}$
$\;-\Bigl{[}P(A_{1}=B_{1}-1)+P(B_{1}=A_{2})$
$\quad+P(A_{2}=B_{2}-1)+P(B_{2}=A_{1}-1)\Bigr{]}.$
To extract randomness, we consider the NPO problem from (10) with a single constraint—namely, the violation of the CGLMP inequality by a specific value. It is straightforward to see that the classical bound is $2$, while an outer bound for the quantum value can be readily computed using the NPA hierarchy, which is approximately $2.91485$. In Figure 3 we present our numerical results. Here, the min-entropy is computed from (25) at NPA level $3$, and the bounds from Theorem II.1 are obtained at NPA level $2$ by interchanging the sum and the infimum, using a fine grid (around 30 points). The numerical computations can be completed in seconds to minutes. As shown in Figure 3, the bounds for randomness extracted via Theorem II.1 are strictly higher when using the CGLMP inequality.
222.22.22.42.42.62.62.82.83300.50.5111.51.5CGLMP valueRandomnessmin-entropy (25)Theorem II.1
(a) min-entropy vs. lower bounds from Theorem II.1 for CGLMP for 2233 scenario
4.84.84.854.854.94.94.954.955500.50.511I3322 valueRandomnessmin-entropy (25)Theorem II.1
(b) min-entropy vs. lower bounds from Theorem II.1 for I3322
Figure 3: Lower bounds for the asymptotically extractable randomness $H(A|E)$ in a device independent scenario are shown in dependence of the violation of the Bell inequalities from [35] and the I3322 Bell inequality. The min-entropy values are obtained by optimizing (25) at NPA level $3$, while the direct bounds are obtained with Theorem II.1 and an appropriate fine grid. The behaviour of the CGLMP inequality is qualitatively very similar to the well-known CHSH case, which we have considered in Figure 2. In comparison to that, the I3322 Bell inequality needs a very high violation $\approx 4.83$ (where 44 is the classical, $\approx 5.0035$ the quantum bound) for extraction of randomness. All results can be achieved on personal devices in the amount of minutes.
IV.2 Randomness from I3322
In this section we consider the generalization of the six-state-protocol originally proposed on qubits in [37]. The idea is to consider three PVM’s per party with two outcomes and to generate key out of one of the three bases. In [37] it is shown that the protocol exhibits slightly improved rates if we consider a device dependent scheme on a qubit. To the best of our knowledge, there are no results for fully device independent lower bounds on the conditional von Neumann entropy in dependence of I3322 violation. We provide in Figure 3 bounds for the asymptotically extractable randomness of the I3322 inequality given by
$I_{3322}=$ $\langle A_{1}B_{3}\rangle+\langle A_{2}B_{3}\rangle+\langle A_{3}B_{1}\rangle$ (29)
$+\langle A_{3}B_{2}\rangle+\langle A_{1}B_{2}\rangle+\langle A_{2}B_{1}\rangle-\langle A_{2}B_{2}\rangle$
$-\langle A_{1}B_{1}\rangle+\langle A_{1}\rangle-\langle A_{2}\rangle+\langle B_{1}\rangle-\langle B_{2}\rangle.$
It is straightforward to see that the classical bound is $4$, while an outer bound for the quantum value can be readily computed using the NPA hierarchy, which is approximately $5.0035$. In Figure 3 we present our numerical results. Here, the min-entropy is computed from (25) at NPA level $3$, and the bounds from Theorem II.1 are obtained at NPA level $2$ by interchanging the sum and the infimum, using a fine grid.
V Test on Experimental Data
In this section we report on the combination of our tools with a set of data measured in a real DI-QKD experiment in lab. For this purpose, we performed an analysis of data taken in a early test run of the DI-QKD experiment [7] showcasing that more measurements actually lead to higher certifiably randomness extraction rates in a realistic environment. In order to gain insight into the asymptotically extractable randomness directly from the observed statistics—rather than first estimating a Bell inequality and then optimizing the conditional von Neumann entropy with respect to its violation—several steps are required. Real experimental data inevitably contain small errors and are therefore only approximately non-signalling. However, at the heart of the NPA hierarchy lies the assumption that we work within the commuting operator framework, which inherently imposes a form of non-signalling condition. As a result, simply applying an $\varepsilon$-threshold to the data may yield a bound corresponding to a not-non-signalling distribution, yet this distribution might still be infeasible at a higher level of the NPA hierarchy. For this reason, when aiming to exploit the full informational content of the distribution, we compute the closest non-signalling distribution to the data and enforce it through equality constraints. Since the conditional von Neumann entropy is a continuous function, this approach does introduce a controllable error in the result, while substantially improving numerical stability.
In this section, we present the numerical results of a 2422-experiment, where we have full access to the distribution $p(a,b\mid x,y)$. Table 2 summarizes the outcomes for one-sided randomness extraction. We compare two scenarios for applying Theorem II.1: one uses fewer grid points, and the other interchanges the sum and infimum but employs a large number of grid points. Additionally, we examine two data-extraction strategies: one that uses only the CHSH value from the statistics (approximately $2.5958$), and another that utilizes the complete distribution. Our results show that, up to a precision of $10^{-4}$, there is no significant difference between using only the CHSH value and using the full statistics. In particular we can compare our results to the theoretical result given by [23] which yields for a CHSH violation of $2.5958$ approximately $0.5759$ bits.
A B C
x values
$x=0$ 0.3883 0 0.3915
$x=1$ 0.4178 0 0.4186
$x=2$ 0.6235 0.6235 0.6156
$x=3$ 0.5888 0.5888 0.5779
y values
$y=0$ 0.6072 0.6069 0.5979
$y=1$ 0.6023 0.6023 0.5921
Table 2: In column A we summarize our certified randomness with $22$ grid points in Theorem II.1 but sum and infimum are interchanged and we have used the full statistics as equality constraints. In B we just incorporated the constraints regarding $x=2,3$ and $y=0,1$ which would correspond to statistics to calculate the CHSH value. In C we do not interchange sum and infimum, used the full statistic but use only 77 grid points. All experiments are done at NPA level 22 and include extra monomials of the form $M_{a|x}N_{b|y}P_{k}^{(a)}$. The results are rounded to a precision $\equiv 10^{-4}$.
As a last contribution we report on two-sided randomness extraction whereby we used $22$ grid points in Theorem II.1, in which the sum and infimum operations were interchanged, and full statistical constraints were employed at NPA level $2$ with extra monomials of the form $M_{a|x}N_{b|y}P_{k}^{(a)}$. For the measurement settings $x=1$ and $y=0$, the global randomness value was found to be approximately $0.8520$. Comparing this with the results in Figure 2, yields that it approximately fits exactly what we would expect from CHSH inequality.
VI Comparison to the method of Brown, Fawzi, and Fawzi
In order to demonstrate the potentials our method offers, we will compare it to the method of Brown, Fawzi, and Fawzi (BFF) [22]. Conceptually there are some parallels, both methods build on lower bounding the von Neumann entropy by upper bounding a relative entropy and both methods use integral representation in order to get approximations. The method of Tan, Schwonnek et. al. [21] follows a different ansatz, in this work we will leave a comparison to this method aside. In the device-independent setting, the definition of the conditional von Neumann entropy becomes significantly more challenging, since its standard formulation via matrix logarithms cannot be used. At present, there are essentially two different integral representations of the relative entropy and, in the case of a classical conditioning system, the conditional von Neumann entropy: Kosaki’s formula [38] and Frenkel’s formula [18]. Both formulations provide a way to define the conditional von Neumann entropy in a device-independent setting.
Kosaki’s formula arises from an interpolation result and can be generalized to all operator monotone functions, whereas Frenkel’s formula belongs to the category of $f$-divergences. Since both techniques rely on integration, they can be compared fairly in terms of their efficiency with respect to:
  1. 1.
    the quality of the lower bound,
  2. 2.
    the number of nodes required,
  3. 3.
    runtime in a competitive example, and
  4. 4.
    the size of the moment matrices.
We provide evidence for the comparison with the help of the following examples. The computational experiments were performed on a system equipped with a 13th Gen Intel® Core™ i5-13600KF processor and 128 GB of DDR4 memory operating at a configured speed of 4000 MT/s (maximum rated speed 4800 MT/s).
VI.1 I3322 comparison
4.64.64.84.85500.50.511Bell ScoreEntropy BoundBFF Method, 8 notesThis work, 8 notesThis work, 30 notes
Figure 4: Comparison of entropy bounds obtained using the BFF method and the from this work as a function of the score. On this instance 8 notes is already significanly challening in runtime $\sim 10$s per point. As comparison our method with also 8 notes needs only $\sim 0.5$s (see Figure 5) and gives higher entropies. We also benchmark our method with 30 notes. The runtime is on the order of $\sim 2$s and we can see that the entropy bound further increases.
We investigate randomness extraction from correlations violating the $I_{3322}$ Bell inequality using two approaches: the Brown–Fawzi–Fawzi (BFF) method with relaxations (1), (3), and (4) from [22, Rem. 2.6], and our proposed technique, for which the analogues of [22, Rem. 2.6 (3),(4)] can also be applied. In both methods, we employ $8$ nodes in the quadrature formula for the integration and work at NPA level 2, supplemented with the minimal number of level 3 moment matrix elements required to ensure that the objective functional is well defined. This setup results in moment matrices of size $110\times 110$ for the BFF method and only $62\times 62$ for our technique.
In Figure 4, we present the lower bounds obtained from Theorem II.2 and BFF for both methods as a function of the Bell score. We conclude that they are on a comparable scale. In Figure 5 we provide a runtime comparison: the BFF method requires between approximately $3.58$ s and $18.70$ s per instance over the considered score range, while our method completes in 0.430.430.510.51 s per instance. This demonstrates that, while both methods produce comparable bounds, our approach achieves a substantial speed-up—often exceeding an order of magnitude—due to the significantly smaller moment matrix size resulting from the simpler objective function and the fact that we can assume the operators in Eve’s system to be projective.
4.64.64.84.855055101015152020Bell ScoreRuntime (s)BFF MethodThis work
Figure 5: Runtime comparison of the BFF method and our method (KS) for randomness extraction from the $I_{3322}$ Bell inequality, using 88 quadrature nodes and NPA level 2 with minimal level 3 extensions.
VI.2 CGLMP comparison
222.22.22.42.42.62.62.82.83300.50.5111.51.5Bell ScoreEntropy BoundBFF MethodThis work
Figure 6: Comparison of entropy bounds obtained using the BFF method and the method from this work for the CGLMP inequality as a function of the score.
We investigate randomness extraction from correlations violating the CGLMP Bell inequality using two approaches: the Brown–Fawzi–Fawzi (BFF) method with relaxations (1), (3), and (4) from [22, Rem. 2.6], and our proposed technique, for which the analogues of [22, Rem. 2.6 (3),(4)] can also be applied. In both methods, we employ $8$ nodes in the quadrature formula for the integration and work at NPA level 2, supplemented with the minimal number of level 3 moment matrix elements required to ensure that the objective functional is well defined. This setup results in moment matrices of size $230\times 230$ for the BFF method and only $122\times 122$ for our technique.
In Figure 6, we present the lower bounds obtained from Theorem II.2 and BFF for both methods as a function of the Bell score, demonstrating comparable scales. In Figure 7 we provide a runtime comparison: across the considered score range, the BFF method requires approximately 69.0569.05242.81242.81 s per instance, while our method completes in only 3.083.083.673.67 s per instance. This illustrates a substantial speed-up of our approach, attributable to the significantly smaller moment matrices resulting from a simpler objective functional and the ability to assume projective operators on Eve’s system.
222.22.22.42.42.62.62.82.8330100100200200Bell ScoreRuntime (s)BFF MethodThis work
Figure 7: Runtime comparison of the BFF method and our method (KS) for randomness extraction from the CGLMP inequality, using 88 quadrature nodes and NPA level 2 with minimal level 3 extensions.
VII Summary
@@
This work focuses on optimizing the conditional von-Neumann entropy within a device-independent framework. From a broader perspective, we ask how much randomness can be extracted from two spatially separated, hypothetical “black boxes” located in Alice’s and Bob’s respective labs, where the outcomes follow the statistical laws of quantum theory. Instead of specifying a concrete quantum experiment represented by a Hilbert space $\mathcal{H}$ and a trace-class operator $\rho$ on $\mathcal{H}$, we rely on the achievable statistics defined by general quantum theory. The core principle underlying the presence of genuine randomness here is the monogamy of correlations, which restricts information sharing in non-signalling theories as demonstrated in [11]. This principle remains applicable even with only limited knowledge of the output statistics in (1). Additionally, bounding the capabilities of Alice, Bob, and Eve to the framework of quantum theory is significant not only for applications but also for advancing our understanding of quantum principles. For instance, [39] exploits the gap between general non-signalling theories and quantum theory, highlighting this difference as a powerful tool.
@@
Our approach serves as a versatile framework for deriving lower bounds on the conditional von-Neumann entropy. The main technical contributions of our work include the application of the integral representation from [18] adapted for this setting, as well as the formulation of a non-commutative polynomial optimization problem, which we solve using the NPA hierarchy [19]. This approach bridges the gap between two established methods: those in [22] and [21]. The technique from [22] has become standard in security proofs for device-independent quantum key distribution (DIQKD) and has since been extended to Petz-Rényi divergences [40]. Compared to [22], our method requires fewer NPA variables, as our projections $P_{k}^{(a)}$ in Theorem II.2 reduce complexity relative to the non-Hermitian operators $Z_{a,i}$ in [22]. This reduction streamlines the operator set in the sum-of-squares cone and enables us to impose additional constraints within the kernel of the map $\varphi$ in [30, Fundamental Lemma]. Moreover, our method builds on discretization techniques for integration from [28], while [22] relies on the Gauss-Radau quadrature method from numerical integration theory. However, the advantages of each integration method for the specific computations required remain open for investigation. The method in [21], though distinct from these approaches, recasts the conditional von-Neumann entropy in (2) as
@@
interpreting it as an entropy production. Here, the channel $T_{Q_{A}\to A}$ acts as the measurement channel on Alice’s system. The method in [21] establishes lower bounds through a variational approach, employing the Gibbs variational principle and Golden-Thompson inequalities. A direct comparison of these conceptually related integral-based methods with the entropy production approach would be a valuable avenue for future research. In section III, we present multiple scenarios to illustrate the effectiveness of our method for randomness extraction. These scenarios include both numerical validations and theoretical comparisons to benchmark results from existing studies.
@@
Firstly, we analyze a $2222$ - scenario (two measurement settings per party, each with two outcomes) under CHSH constraints, which is a standard setting in quantum information theory for testing Bell inequalities. We consider this scenario for both one-sided and two-sided randomness extraction, where randomness is certified by outcomes that are either private to one party (one-sided) or to both parties (two-sided). Our results in this $2222$ - scenario show strong agreement with the analytical values reported in [11], demonstrating the accuracy and robustness of our approach. Additionally, we explore a more intricate $2322$ - scenario, where one party has three measurement settings while the other has two. This scenario is used to examine global randomness extraction, which involves jointly certifying randomness across all outcomes. Our findings, obtained under the statistical framework of Werner states, indicate that this more complex setup does not offer a substantial qualitative advantage over the $2222$ - scenario. This suggests that increasing the measurement settings in this configuration may not necessarily enhance randomness extraction, a result that merits further investigation. An intriguing direction for future research would be to systematically compare different configurations - such as the $2222$, $2322$ and $3322$ - scenarios (where the latter allows each party three measurement settings with two outcomes) - to determine how these variations impact randomness generation under different assumptions of honest implementation or even a random key basis [26]. This could uncover new insights into the optimal design of randomness extraction protocols. Moreover, it would be valuable to examine randomness extraction in more complex non-local games, such as the magic square game discussed in [39]. These games offer alternative frameworks for randomness certification that might provide unique advantages in terms of security and efficiency in quantum protocols.
VIII Acknowledgements
GK thanks Mario Berta, Omar Fawzi, Ian George, Edwin Lobo, Stefano Pironio and Lewis Wooltorton for fruitful discussions. GK thanks Marco Tomamichel for hosting him at CQT Singapore in fall 2024. The author would like to thank Rutvij Bhavsar for insightful discussions and for providing access to the data from [32]. GK acknowledges support from the Excellence Cluster - Matter and Light for Quantum Computing (ML4Q). GK acknowledges funding by the European Research Council (ERC Grant Agreement No. 948139). R.S. is supported by the DFG under Germany’s Excellence Strategy - EXC-2123 QuantumFrontiers - 390837967 and SFB 1227 (DQ-mat) , the Quantum Valley Lower Saxony, and the BMBF projects ATIQ, SEQUIN, QuBRA and CBQD. The authors thank Harald Weinfurter and his research group for measuring the test data that was used for benchmarking our methods.